Circular No. 8222

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2003 T2 (LINEAR) An apparently asteroidal object reported by LINEAR (discovery ob-

servation below), and posted on the NEO Confirmation Page, has been found to show cometary appearance by several CCD astrometric observers, including R. Apitzsch (Wildberg, Germany, 0.24-m reflector; Oct. 14.0 UT, diffuse coma of diameter ∼ 10′′), J. E. McGaha (Tucson, AZ, 0.62-m reflector; Oct. 14.4, bright compact coma of diameter 4′′ with a fainter outer coma of diameter 10′′ and a broad tail 20′′ long in p.a. 30^{o}), G. R. Jones (Tucson, AZ, 0.32-m reflector; Oct. 14.4, coma diameter ∼ 6′′ and a possible tail at p.a. 35^{o}), P. R. Holvorcem and M. Schwartz (Tenagra 0.81-m reflector at Nogales, AZ; Oct. 14.5, co-addition of three 120-s exposures shows a coma of diameter ∼ 15′′ and total mag 15.1–15.6), and A. Knöfel and T. Payer (Essen, Germany, 0.32-m reflector; Oct. 14.9, short tail).

2003 UT  α_{2000}  δ_{2000}  Mag.
Oct. 13.44392  9^{h}52^{m}52.^{s}92  +79^{o}54′21′′.6  18.3

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2003-T70.

T = 2003 Nov. 14.830 TT  ω = 152.^{o}941
Ω = 238.427 }2000.0
q = 1.77599 AU  i = 87.472

NOVAE IN M31 Further to IAUC 8165, two apparent novae in M31 have been discov-

ered by K. Hornoch (Sept. 30.835 UT, R = 17.0, α = 0^{h}42^{m}46^{s}.72, δ = +41^{o}19′46′′.7, equinox 2000.0, or 26′′ east and 219′′ north of the galaxy center) and by M. Fiaschi, F. Di Mille, and D. Tiveron (Oct. 6.908, [Hα] = 16.7, α = 0^{h}42^{m}03^{s}.11, δ = +41^{o}14′50′′.6, or 465′′ west and 78′′ south of the galaxy center). Additional R-band magnitudes from Hornoch and [Hα] magnitudes by Fiaschi et al. (F) of the object found by Hornoch: Sept. 26.812, [18.6; 30.835, 17.0; Oct. 1.990, 17.2; 6.908, 16.6 (F); 9.908, 15.7 (F); 12.753, 17.3. Additional magnitudes of the second object: Aug. 22.980 UT, [18.8; 23.971, 18.1; 24.869, 17.4; 26.930, 17.6; 26.976, 16.2 (F); 28.908, 17.8; Sept. 6.933, 18.0; 17.810, 18.3; 21.822, 18.6; 24.928, 18.6; 30.825, [18.6; Oct. 9.908, 16.5 (F). The apparent nova found by Rector (cf. IAUC 8210) also appears at mag 16.6 on an Hα image taken by Fiaschi et al.

2003 October 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green